OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

1. find the slope first i.e. m = y2-y1/x2-x1

OpenStudy (anonymous):

okay after that?

OpenStudy (anonymous):

2. use the equation y-y1 = m(x-x1)

OpenStudy (anonymous):

so for the first part you should have got (-4-0)/(4-0) = -4/4 = -1

OpenStudy (anonymous):

for the second part we let (x1, y1) = (4,-4) so our equation becomes: y-(-4) = -1(x-4) y+4 = -x + 4 y=-x This works equally well if you chose (0,0) as(x1,y1)